if 60% of 75% of a number is 270, then find 40% of that number?
Answer by ptaylor(2198)

Let x=that number
40% of that number is 0.40x
75% of that number is 0.75x
60% of 75% of that number is 0.60(0.75x)=0.45x and we are told that this equals 270
so:
0.45x=270
x=600
0.40x=0.40*600=240
CK
75% of 600 =0.75*600=450
60% of 450=0.60*450=270
40% of 600=240
Hope this helps---ptaylor